# BECS: A Privacy-Preserving Computing Sharing Mechanism for 6G Computing Power Network

### v2

@@ III-B Communication Model
The NOMA-based 6G networks, regarded as the most promising future multi-access technology, is considered in BECS. It allows multiple users to transmit on the same frequency band, thereby enhancing spectrum resource utilization [26]. This model primarily addresses the communication latency in wireless communications, thus overlooking the latency in wired connections. Let $g_{u}$ denote the instantaneous channel gain from $U_{u}$ to the offloading computing device. Without loss of generality, the channel gains for all $U$ devices are ordered in descending sequence: $g_{1}<\cdots<g_{u}<\cdots<g_{U}$. Using successive interference cancellation and decoding for the user with the highest received power, the data transmission rate for $U_{u}$ during the computing offloading is calculated via Shannon’s theorem as follows:
$R_{u}=B\log_{2}\left(1+\frac{p_{u}g_{u}}{\sigma_{0}^{2}+\sum_{n=u+1}^{N}p_{n}g_{n}}\right).$ (1)
@@ III-C Computing Model
In the considered three layers computing sharing structure, the computing tasks of a CRR can be executed in any device of CRP. However, computing devices vary in terms of task transfer and processing capabilities. This intense competition for computing resources at each layer, inevitable in 6G PCN with numerous CRRs and CRPs, necessitates balancing the task load across the system’s layers. Generally, with limited computing resources at each layer, tasks may have to wait for an available processor. Therefore, the M/M/c model [27], which describes task offloading as a Poisson process with an average arrival rate of $\lambda_{t}$, can be used to model task processing delays. Based on the superposition property of the Poisson process and the offloading interactions among the three layers of computing resources, the average task arrival rate for each layer can be calculated as
$\lambda_{d}=\begin{cases}\sum_{t=1}^{T}\lambda_{t}\times p_{d^{*}\to u,t}&\text{if }d=u\\ \sum_{t=1}^{T}\sum_{d^{*}\in\{u,e\}}\lambda_{t}\times p_{d^{*}\to e,t}&\text{if }d=e\\ \sum_{t=1}^{T}\sum_{d^{*}\in\{u,e,c\}}\lambda_{t}\times p_{d^{*}\to c,t}&\text{if }d=c\end{cases}.$ (4)
@@ III-D Service Model
1) Computing occupancy: As a direct indicator of computing resource utilization, it is quantified by the number of devices participating in computing sharing while maintaining the provision of all network services. It aims to enhance the overall breadth of resource engagement across the system, which can be defined as follows:
@@
2) Privacy entropy: More applicable computing offloading can be achieved by constructing three layers computing structures. However, the decentralization of computing devices introduces privacy leakage risks during task offloading. Privacy entropy, a method of quantifying privacy security, is used to measure the security of information transmission [29]. Higher privacy entropy indicates more disordered data, preventing attackers from making inferences and thereby ensuring secure information transmission.
The types of offloaded data vary according to the computing tasks. Thus, in BECS, we set the relationship between $T_{t}$ and $p_{d^{\ast}\to d,t}$ to be a one-to-one correspondence, and the privacy entropy of $T_{t}$ in $D_{d}$ can be calculated as
@@
4) Sharing revenue: A suitable revenue can encourage the participation of computing devices in computing sharing, thereby enhancing the overall resource utilization. The logarithmic utility function is used to quantify the sharing revenue of $D_{d}$, which can be calculated as
@@ III-E Premise Assumptions
1) Computing tasks can be offloaded from the outer to the inner layer as depicted in Fig. 1, or within the same layers, provided that the computing capacity of CRP exceeds that of CRR.
2) Each user device is equipped with a single antenna, facilitating real-time communication with the 6GBS.
3) Edge and cloud computing devices are typically connected to the 6GBS through high-speed, reliable wired connections, such as optical fibers. The latency of these connections is significantly lower than that of wireless communication between user devices and the 6GBS, and can be considered negligible.
4) Based on the permissioned blockchain, user registration information is visible only to 6GBSs within the blockchain, whereas the computing resources status is visible to all users.
5) All users securely deliver their keys through a secure channel.
@@ IV-A MOOP of Computing Allocation
Based on the aforementioned model, total time and energy consumption can serve as metrics for computing cost and green 6G, which can be calculated as follows:
@@
and
@@
In addition, to ensure the user’s service experience, the computing resource utilization the average privacy entropy of the computing tasks, the load balancing of the computing devices, and the total sharing revenue of CRP are considered, which can be calculated as follows:
@@
and
@@
Assigning tasks to free computing devices can effectively improve the resource utilization of the entire system. However, focusing solely on utilization improvement is unscientific. Therefore, it is crucial to analyze the trade-offs among these intertwined and conflicting objectives. In BECS, our objectives include minimizing time and energy consumption, optimizing load balancing to maximize resource utilization, ensuring user experience, and simultaneously maximizing privacy entropy and CRP revenue. The MOOP is formulated as:
$\max\left\{H_{ave},R_{tot},O_{ave}\right\},$ (19)
@@
$\zeta_{u}<\zeta_{e}<\zeta_{c},\zeta_{d^{*}}<\zeta_{d},$ (25)
$\xi_{d,t}\leq s_{d}$ (26)
Constraint (21) states that the actual task processing rate of the computing device must not exceed its service rate. Constraint (22) restricts the maximum transmit power of the user device. Constraint (23) guarantees that the task execution time does not exceed its deadline. Constraint (24) specifies that the computing energy consumption should not exceed the device’s maximum available energy. Constraint (25) differentiates the computing capacities across three layers and specifies task offloading from devices with lower capacity to those with higher capacity. Constraint (26) stipulates that the computing demands of tasks offloaded to executing devices must not exceed their capacities.
Figure 2: Encoding of computing resources.
@@ IV-B Kernel Distance-based Dominance Relation
NSGA-III, an excellent evolutionary algorithm, can achieve fast global searches with quality assurance, preventing MOOP from settling into local optimality. Additionally, it tackles high-dimensional problems by preserving population diversity through uniformly distributed reference points [31]. In practical MOOP, maximizing population diversity while ensuring convergence is one of the best ways to quickly and effectively obtain the optimal solution. Consequently, NSGA-III-KDR is proposed to balance the aforementioned five objectives.
The performance of Pareto dominance relation-based algorithms often exhibits serious dimensionality implications due to dominance resistance in MOOP with more than three objectives [32]. To address this, Tian et al. [13] proposed a strengthened dominance relation (SDR) to enhance NSGA-II. Considering that NSGA-III builds upon NSGA-II by introducing reference points, we propose a kernel distance-based dominance relation (KDR) in NSGA-III-KDR. It builds on the theoretical foundation of SDR to replace the original Pareto dominance relation in NSGA-III. Specifically, if solution $X_{1}$ dominates solution $X_{2}$ in KDR, then (27) is satisfied.
@@ IV-C Encoding of Computing Resources
In BECS, the occupancy status of computing devices is encoded as genes, while the computing resources involved in sharing are encoded as chromosomes, constituting the entire CRP for evolution. As illustrated in Fig. 2, a gene value of 0 indicates a free computing device, whereas a value of 1 signifies that this device is occupied. This method allows BECS to integrate various types of computing resources, thereby building a generalized computing sharing platform. Since the occupancy status of computing devices directly correlates with gene encoding, and chromosomes relate to computing allocation strategies, dynamic and fine-grained updates of computing resources are enabled.
@@
0:  The initialized population $X^{1}$, the total number of iterations $TI$, and the population size $PS$
@@
IV-D Population Evolution
When the CRR requests computing resources, the system executes the computing allocation scheme based on NSGA-III-KDR to match the optimal CRP. NSGA-III-KDR basically follows the algorithmic framework of NSGA-III, as shown in Algorithm 1. During the evolution process, the chromosomes in the initial population $X^{1}$ generate entirely new chromosomes (computing allocation strategies) through crossover and mutation operations. The crossover operation enhances chromosome diversity, while the mutation operation, under specific conditions, modifies individual genes to seek those with higher adaptability. In each iteration, the population generated by crossover and mutation merges with the original, then executing a Non-KDR-dominated sort that replaces the non-dominated sort in NSGA-III, thereby effectively enhancing the diversity of the computing resource population. Subsequently, by associating chromosomes with reference points in the hyper-plane and executing the Niche-Preservation Operation, the next generation population can be generated. After $TI$ iterations of NSGA-III-KDR, the final population $X^{TI}$, containing the ultimate computing allocation strategy, is obtained.
@@ IV-E Optimum Selection
The final population $X^{TI}$ comprises a set of the best feasible solutions for computing allocation, termed Pareto solutions. However, in practical computing allocation problems, the CRR needs to match only a specific CRP. Therefore, the Technique for Order Preference by Similarity to Ideal Solution (TOPSIS), in conjunction with entropy weighting, is utilized to evaluate the most optimal solution among many Pareto solutions. As a method approximating the ideal solution, TOPSIS rapidly identifies the optimal solution by ranking all solutions according to their distance from the positive (negative) ideal solution. Additionally, entropy weighting helps eliminate the arbitrariness of subjectively determined weights. This combination enables an objective and fair determination of the optimal computing allocation from $X^{TI}$. The specific algorithm is detailed in [35].
By executing NSGA-III-KDR in BECS, CRP can set $\psi_{d}$ more reasonably by analyzing the optimal solution. Simultaneously, CRR can select the most appropriate CRP for task offloading based on the optimal solution. Consequently, both parties engaged in computing sharing achieve reliable, dynamic, and secure computing allocation and trading.
@@ V-A System Initialization
In this phase, upon inputting the security parameter $\lambda$, the system administrator selects a secure elliptic curve $\mathcal{E}:y^{2}=x^{3}+ax+b\mod p$. The group $\mathcal{G}$ is an additive elliptic curve group with order $q$ and generator $g$ defined over $\mathcal{E}$, where $p$ and $q$ are two large prime numbers. Subsequently, 6GBS $B_{b}$ chooses a random number $b\in Z_{q}^{*}$ as its private keys and caculates the public keys as $bg$. Additionally, it selects a collision-resistant hash function ():$\mathcal{H()}:\{0,1\}^{*}\to\{0,1\}^{l}$ and makes $\{\mathcal{E},\mathcal{G},x,p,q,\mathcal{H()}\}$ publicly available in blockchain.

### v3

@@ III-B Communication Model
In BECS, a Non-Orthogonal Multiple Access (NOMA) based communication model is considered for the 6G uplink, which allows multiple user devices to transmit data to the 6GBS over the same frequency band, thereby significantly enhancing spectral efficiency [20, 78, 79]. Successful decoding in NOMA hinges on the Successive Interference Cancellation (SIC) technique at the receiver [88, 59]. This requires the signals to be decoded sequentially, which is predicated on a known channel gain order. Specifically, the receiver decodes signals in descending order of their channel gains; crucially, the power allocation is constrained to guarantee a minimum data rate for weak users, ensuring fairness and satisfying their Quality of Service lower bound.
Without loss of generality, the users are indexed such that their channel gains are in descending order, i.e., $g_{1}>\cdots>g_{u}>\cdots>g_{U}$. Consequently, when decoding the signal for user $U_{u}$, signals from users with stronger channel conditions ($U_{1},\dots,U_{u-1}$) have already been cancelled. The signals from users with weaker channel conditions ($U_{u+1},\dots,U_{U}$) constitute the co-channel interference. The resulting data transmission rate for $U_{u}$ is therefore calculated via Shannon’s theorem as follows:
$R_{u}=B\log_{2}\left(1+\frac{p_{u}g_{u}}{\sigma_{0}^{2}+\sum_{n=u+1}^{U}p_{n}g_{n}}\right).$ (1)
@@ III-C Computing Model
In the considered three layers computing resource sharing structure, the computing tasks of a CRR can be executed in any device of CRP. However, computing devices vary in terms of task transfer and processing capabilities. This intense competition for computing resources at each layer, inevitable in 6G CPN with numerous CRRs and CRPs, necessitates balancing the task load across the system’s layers. Generally, with limited computing resources at each layer, tasks may have to wait for an available processor. Therefore, the M/M/c model [21, 81, 82], which describes task offloading as a Poisson process with an average arrival rate of $\lambda_{t}$, can be used to model task processing delays. Based on the superposition property of the Poisson process and the offloading interactions among the three layers of computing resources, the average task arrival rate for each layer can be calculated as
$\lambda_{d}=\begin{cases}\sum_{t=1}^{T}(\lambda_{t}\times p_{d^{*}\to u,t})&\text{if }d=u\\ \sum_{t=1}^{T}\sum_{d^{*}\in\{u,e\}}(\lambda_{t}\times p_{d^{*}\to e,t})&\text{if }d=e\\ \sum_{t=1}^{T}(\sum_{d^{*}\in\{u,e,c\}}\lambda_{t}\times p_{d^{*}\to c,t})&\text{if }d=c\end{cases}.$ (4)
@@ III-D Service Model
1) Computing occupancy: As a direct indicator of computing resource utilization, it is quantified by the number of devices participating in computing resource sharing while maintaining the provision of all network services. It aims to enhance the overall breadth of resource engagement across the system, which can be defined as follows:
@@
2) Privacy entropy: Beyond the privacy of task content, the decentralization of computing devices in the 6G CPN introduces significant privacy risks stemming from the offloading patterns themselves. Deterministic offloading strategies, while optimal in terms of performance, generate predictable behavioral patterns. As these decisions are often correlated with the user’s wireless channel state—and consequently their location—a curious adversary can analyze offloading metadata (e.g., task size and frequency) to infer sensitive information. To mitigate such statistical inference attacks, privacy entropy is employed to quantify the unpredictability of the task allocation process [23, 83, 84]. A higher entropy indicates a more disordered allocation pattern, which makes it computationally difficult for an adversary to perform accurate inferences, thereby enhancing user privacy. The types of offloaded data vary according to the computing tasks. Thus, in BECS, we set the relationship between $T_{t}$ and $p_{d^{\ast}\to d,t}$ to be a one-to-one correspondence, and the privacy entropy of $T_{t}$ in $D_{d}$ can be calculated as
@@
4) Sharing revenue: A suitable revenue can encourage the participation of computing devices in computing resource sharing, thereby enhancing the overall resource utilization. The logarithmic utility function is used to quantify the sharing revenue of $D_{d}$, which can be calculated as
@@ III-E Premise Assumptions
1) The proposed computing resource sharing operates within a metropolitan area, where edge and cloud nodes constitute a well-connected computing infrastructure, linked by a high-capacity optical backbone.
2) Computing tasks can be offloaded from the outer to the inner layer as depicted in Fig. 1, or within the same layers, provided that the computing capacity of CRP exceeds that of CRR.
3) Each user device is equipped with a single antenna, facilitating real-time communication with the 6GBS.
4) The communication latency on the optical backbone is assumed to be negligible in comparison to the dominant delays introduced by wireless access and task computing [61, 89].
5) Based on the permissioned blockchain, user registration information is visible only to 6GBSs within the blockchain, whereas the status of computing resources is visible to all users.
6) All users securely deliver their keys through a secure channel.
@@ IV-A MOOP of Computing Allocation
The core of our computing allocation mechanism is formulated as a MOOP. The goal is to intelligently assign computing tasks to available resources in the 6G CPN, balancing multiple, often conflicting, performance objectives. These objectives span from system efficiency and cost to user-centric metrics like privacy and service quality.
To comprehensively address the critical requirements of the 6G CPN ecosystem, including Service Quality and Green Sustainability, System Stability and Efficiency, User Privacy, and Economic Viability, we identify six representative objectives that capture the fundamental trade-offs inherent in decentralized resource sharing. First, the system cost objectives are defined, which include the total time consumption ($L_{tot}$) and total energy consumption ($E_{tot}$). These serve as key metrics for evaluating computing efficiency and energy sustainability in green 6G networks.
Total Time Consumption ($L_{tot}$): As formulated in (13), this metric captures the overall latency. It is the sum of the computation time ($L^{com}$) across all three layers and the wireless transmission time ($L^{tra}$) incurred when user devices offload tasks.
@@
Total Energy Consumption ($E_{tot}$): Similarly, as shown in (14), this objective accounts for the total energy spent. It comprises the energy for task execution ($E^{com}$) on the designated devices and the energy for wireless transmission ($E^{tra}$) from user devices.
@@
Next, a set of objectives is defined to address key aspects of system performance, including resource utilization ($O_{ave}$), privacy entropy ($H_{ave}$), load balancing ($B_{tot}$), and CRP revenue ($R_{tot}$).
Average Resource Utilization ($O_{ave}$): As formulated in (15), the average resource utilization quantifies the extent of resource engagement by representing the ratio of occupied devices ($O_{tot}$) to the total number of devices ($D$).
@@
Average Privacy Entropy ($H_{ave}$): As shown in (16), the average privacy entropy serves as an indicator of user privacy protection. A higher individual entropy ($H_{d,t}$) reflects a more disordered and less predictable task allocation pattern, thereby increasing the difficulty for adversaries to infer sensitive information.
@@
Total Load Balancing ($B_{tot}$): As expressed in (17), the total load balancing metric is computed by aggregating the individual load balance indicators $B_{d}$ across all devices. The underlying indicator $B_{d}$, defined in (11), corresponds to the standard deviation of workloads, where a smaller value indicates a more balanced load distribution and contributes to the mitigation of computational bottlenecks.
@@
Total CRP Revenue ($R_{tot}$): As given in (18), the total CRP revenue is obtained by aggregating the individual revenues ($R_{d}$) of all resource-providing devices. The individual revenue function $R_{d}$, detailed in (12), is formulated as a logarithmic utility to model economic incentives, capturing diminishing returns and promoting broad participation.
@@
Assigning tasks to free computing devices can effectively improve resource utilization. However, focusing solely on utilization improvement is insufficient, as the objectives within the 6G CPN are inherently conflicting. While individual metrics such as time and energy consumption are well-established [2, 3, 7], the novelty of this study lies in holistically modeling, for the first time, the complex interplay among performance, efficiency, security, and economy. For instance, minimizing time and energy consumption often leads to workload centralization, which directly conflicts with the goal of equitable load balancing. Similarly, maximizing privacy entropy may necessitate non-optimal routing that degrades performance, whereas a strategy aimed at maximizing provider revenue may create economic tensions that impact overall resource occupancy. This landscape of competing demands necessitates a solution that seeks the best possible compromise. It is worth noting that although other performance indicators (e.g., jitter or packet loss) exist, they are inherently correlated with the selected metrics (e.g., latency) or are less critical under the assumption of a high-capacity optical backbone. Therefore, these six objectives are identified as the minimally sufficient set for characterizing the system trade-offs without introducing redundant dimensions that could unnecessarily hinder algorithmic convergence. Thus, to scientifically balance these competing objectives, the MOOP is formulated as follows:
$\max\left\{O_{ave},H_{ave},R_{tot}\right\},$ (19)
@@
$\varsigma_{u}<\varsigma_{e}<\varsigma_{c},\varsigma_{d^{*}}<\varsigma_{d},$ (25)
$\lambda_{d}\xi_{d}\leq\varsigma_{d}$ (26)
Constraint (21) states that the actual task processing rate of the computing device must not exceed its service rate. Constraint (22) restricts the maximum transmit power of the user device. Constraint (23) guarantees that the task execution time does not exceed its deadline. Constraint (24) specifies that the computing energy consumption should not exceed the device’s maximum available energy. Constraint (25) differentiates the computing capacities across three layers and specifies task offloading from devices with lower capacity to those with higher capacity. Constraint (26) mandates that the average rate of incoming computing workload to a device must not exceed its processing capacity.
@@ IV-B Kernel Distance-based Dominance Relation
NSGA-III, a state-of-the-art evolutionary algorithm, can perform fast global searches with quality guarantees, effectively preventing the MOOP from converging to local optima. Additionally, it addresses high-dimensional optimization problems by maintaining population diversity through uniformly distributed reference points [25]. However, as evidenced by the complexity of the proposed practical six-objective MOOP, even advanced algorithms encounter significant challenges. The performance of Pareto dominance-based algorithms often suffers from severe dimensionality effects caused by dominance resistance when handling MOOPs with more than three objectives [26]. This phenomenon substantially weakens selection pressure, making it difficult for the algorithm to distinguish among a large number of mutually non-dominated solutions, which is an essential limitation when a diverse set of operational strategies is required for the 6G CPN.
To overcome this specific bottleneck within the BECS framework, a more powerful dominance relation is required. NSGA-III-KDR is inspired by the concept of a strengthened dominance relation (SDR), which has been demonstrated to be effective in NSGA-II [27]. Given that the core diversity-preserving mechanism of NSGA-III relies on reference points, a distance-based metric is conceptually a natural choice for enhancing its selection process. Accordingly, we propose a kernel distance-based dominance relation (KDR) in NSGA-III-KDR. It builds upon the theoretical foundation of SDR to replace the original Pareto dominance relation in NSGA-III. This enables the algorithm to evaluate solutions not only based on dominance but also in terms of their proximity to the ideal point, thereby ensuring stronger selection pressure for addressing the complex computing sharing in 6G CPN. Specifically, if solution $X_{1}$ dominates solution $X_{2}$ in KDR, then (27) is satisfied.
@@ IV-C Encoding of Computing Resources
In BECS, the occupancy status of computing devices is encoded as genes, while the computing resources involved in sharing are encoded as chromosomes, constituting the entire CRP for evolution. As illustrated in Fig. 2, a gene value of 0 indicates a free computing device, whereas a value of 1 signifies that this device is occupied. This method allows BECS to integrate various types of computing resources, thereby building a generalized computing resource sharing platform. Since the occupancy status of computing devices directly correlates with gene encoding, and chromosomes relate to computing allocation strategies, dynamic and fine-grained updates of computing resources are enabled.
Figure 2: Encoding of computing resources.
IV-D Population Evolution
To adapt to the dynamic 6G CPN environment, in which task arrivals and resource availability fluctuate over time, BECS operates not as a one-time process but as an adaptive control loop. When a CRR requests computing resources in the network, the computing resource sharing mechanism is triggered. Subsequently, the system executes the NSGA-III-KDR-based computing allocation scheme to match the optimal CRP according to the current network state. NSGA-III-KDR essentially follows the algorithmic framework of NSGA-III, as shown in Algorithm 1. During the evolution process, the chromosomes in the initial population $X^{1}$ generate entirely new chromosomes (computing allocation strategies) through crossover and mutation operations. The crossover operation enhances chromosome diversity, while the mutation operation, under specific conditions, modifies individual genes to seek those with higher adaptability. In each iteration, the population generated by crossover and mutation merges with the original, then executing a Non-KDR-dominated sort that replaces the non-dominated sort in NSGA-III, thereby effectively enhancing the diversity of the computing resource population. Subsequently, by associating chromosomes with reference points in the hyper-plane and executing the Niche-Preservation Operation, the next generation population can be generated. After $TI$ iterations of NSGA-III-KDR, the final population $X^{TI}$, containing the ultimate computing allocation strategy, is obtained.
@@ IV-D Population Evolution
0: The initialized population $X^{1}$, the total number of iterations TITI, and the population size PSPS
@@ IV-E Optimum Selection
The final population $X^{TI}$ comprises a set of the best feasible solutions for computing allocation, termed the Pareto optimal set. Each solution in this set represents a different trade-off among the conflicting objectives and is optimal in the sense that no objective can be improved without degrading at least one other. However, in practical computing allocation problems, the CRR needs to match only a specific CRP, necessitating a decision-making process to select a single strategy for implementation. Therefore, the Technique for Order Preference by Similarity to Ideal Solution (TOPSIS) [85], in conjunction with entropy weighting, is utilized to select the most suitable solution for a specific requirement from the set of Pareto optimal solutions. As a method approximating the ideal solution, TOPSIS rapidly identifies the optimal solution by ranking all solutions according to their distance from the positive (negative) ideal solution. Additionally, entropy weighting helps eliminate the arbitrariness of subjectively determined weights. This combination enables an objective and fair determination of the final computing allocation strategy from the $X^{TI}$. The specific algorithm is detailed in [30].
By executing NSGA-III-KDR in BECS, CRP can set $\psi_{d}$ more reasonably by analyzing the optimal solution. Simultaneously, CRR can select the most appropriate CRP for task offloading based on the optimal solution. Consequently, both parties engaged in computing resource sharing achieve reliable, dynamic, and secure computing allocation and trading.
@@ V-A System Initialization
In this phase, upon inputting the security parameter $\lambda$, the system administrator selects a secure elliptic curve $\mathcal{E}:y^{2}=x^{3}+ax+b\mod p$. The group $\mathcal{G}$ is an additive elliptic curve group with order $q$ and generator $g$ defined over $\mathcal{E}$, where $p$ and $q$ are two large prime numbers. Subsequently, 6GBS $B_{b}$ chooses a random number $b\in Z_{q}^{*}$ as its private key and calculates the public key as $bg$. Additionally, it selects a collision-resistant hash function $\mathcal{H()}:\{0,1\}^{*}\to\{0,1\}^{l}$ and makes $\{\mathcal{E},\mathcal{G},a,b,p,q,\mathcal{H()}\}$ publicly available in blockchain.